# Conserved charges and asymptotic symmetries of BTZ-like black holes in Einstein-bumblebee gravity

### v1

@@ IV Conserved charges for BTZ-like bumblebee black hole
Mass: By choosing the exact symmetry $\eta_{{}_{M}}=\partial_{t}$ and substituting the surface charge expressions (29)-(III) along with the parametric variations (IV) into Eq.(35), we can obtain the total mass
@@
The surface gravity, temperature and angular velocity on the horizon are given by Compere:2018aar
@@
In our formulation the reference points for the charges above are chosen to vanish when $\mu=j=0$. The total mass, angular momentum and entropy are all affected by the factor $\sqrt{1+s}$, arising due to LV effect. The first law is satisfied perfectly. When $s\rightarrow 0$ , this solution reduces to the exact BTZ black hole solution, and $M=\frac{\mu}{4G}$, $J=\frac{j}{4G}$, $S=\frac{\pi r_{{}_{\mathrm{H}}}}{G}$ (up to factor $\frac{1}{2}$ , which comes from the choice of the gravitational coupling constant $\kappa=4\pi G$ in original black hole solution Ding:2023niy and also in our case). As a consistent check, in the next section we will derive the microscopic entropy of dual CFT by the AdS/CFT correspondence, and find that the Bekenstein-Hawking entropy (41) is in precise agreement with the microscopic entropy of the dual CFT.
V Asymptotic symmetries
@@
V.1 AdS/CFT correspondence
Setting $\mu=0=j$ in the metric and bumblebee field of the BTZ-like black hole (IV), we can get the $\mathrm{AdS}_{3}$ background spacetime
@@ V.1 AdS/CFT correspondence
By using the Brown-Henneaux boundary conditions Strominger:1997eq
@@
$i\left[\xi^{+}_{m},\xi^{-}_{n}\right]=$ $0$ (48)
For the generator $\xi^{+}_{m}$, the Poisson bracket algebra (II) becomes
@@
$i\oint_{S}\bm{k}_{\xi^{-}_{m}}(\delta_{\xi^{-}_{n}}\bar{\Phi},\bar{\Phi})=\frac{1}{12}\frac{3\sqrt{1+s}l}{G}m^{3}\delta_{m+n,0}$ (51)
where $\tilde{L}^{+}_{m}\equiv H_{\xi^{+}_{m}}$, $\tilde{L}^{-}_{m}\equiv H_{\xi^{-}_{m}}$, and $\bar{\Phi}$ is the background (V.1). To contain centerless subalgebra for $m=-1,0,+1$, we shift the zero modes of the charges
@@
By using the Cardy formula Cardy:1986ie ; deBoer:2010ac , and modified by the LV factor $\sqrt{1+s}$, the microscopic entropy of the dual CFT is
@@
The zero modes of $\tilde{L}^{+}_{m}$ and $\tilde{L}^{-}_{n}$ for the BTZ-like black hole (IV) are
@@
V.2 Warped conformal symmetries in near horizon region
In this subsection we will use the Detournay-Smoes-Wutte boundary conditions Detournay:2023zni to study the asymptotic symmetries of the near horizon extremal BTZ-like black hole.
In the extremal case, the BTZ-like metric and the bumblebee field (IV) are
@@ V.2 Warped conformal symmetries in near horizon region
By taking limit $\varepsilon\rightarrow\infty$, we can obtain the near horizon geometry of the extremal BTZ-like metric and bumblebee field
@@
We now impose the Detournay-Smoes-Wutte boundary conditions Detournay:2023zni on the near horizon geometry of the extremal BTZ-like metric and bumblebee field (V.2), which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity Godet:2020xpk , and it is given by a finite coordinate transformation
@@
where $\mathcal{H}(\tau)\equiv\mathcal{F^{\prime\prime}}(\tau)/\mathcal{F^{\prime}}(\tau)$. The near horizon geometry (V.2) is restored when $\mathcal{G^{\prime}}(\tau)=0=\mathcal{H}(\tau)$.
@@
$i\left[j_{m},j_{n}\right]=$ $0$ (66)
To obtain finite and non-zero charges (II), we take $\partial{\Sigma}$ to be the circle S of constant $(r,\varphi)$ and take the limit $r\rightarrow\infty$. By taking variations of the metric and bumblebee field
@@
the corresponding charge variation (II) associated with $\xi$ can be given by
[…]
which is integrable. Under the background (V.2), the charges are given by
[…]
The Poisson bracket algebra (II) of $\hat{L}_{n}$ and $\hat{J}_{n}$ is a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, the symmetry algebra of a WCFT,
@@
VI conclusions and remarks
In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole in Einstein-bumblebee theory of gravity. We explicitly computed the black hole mass, angular momentum and entropy, demonstrating that these quantities are influenced by LV parameter. Additionally, we derived the first law of black hole thermodynamics within this framework. As an asymptotical AdS spacetime, we studied the AdS/CFT correspondence and obtained the asymptotic charge algebra, which are two copies of the Virasoro algebra with non-trivial central charges. By applying the Cardy formula, we computed the microscopic entropy of the dual CFT, and found it is in exact agreement with the Bekenstein-Hawking entropy calculated by SPSM. Furthermore, in the near horizon region, we used the Detournay-Smoes-Wutte boundary conditions to study the near horizon geometry of extremal BTZ-like black hole. We obtained a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of a WCFT.
Evidently, the conserved charges and central charges are affected by the LV parameter. We speculate that a similar modification arises in other bumblebee black hole solutions Casana:2017jkc ; Maluf:2020kgf ; Jha:2020pvk ; Filho:2022yrk ; Xu:2022frb ; Liu:2024axg , and also affect the thermodynamic properties. Notably, in the near horizon geometry, the central charge $\hat{c}$ of the Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra is found to be identical to the central charge $c$ of the Virasoro algebra in AdS/CFT correspondence. This result suggests a potential dual holographic description linking extreme black hole to WCFT, providing an alternative means to derive the black hole entropy by this duality.
Acknowledgements.

### v2

@@ 4 Conserved charges for BTZ-like bumblebee black hole
Mass: By choosing the exact symmetry $\eta_{{}_{M}}=\partial_{t}$ and substituting the surface charge expressions (29)-(3) along with the parametric variations (4) into Eq.(35), we can obtain the total mass
@@
The surface gravity, temperature and angular velocity on the horizon are given by [11]
@@
In our formulation the reference points for the charges above are chosen to vanish when $\mu=j=0$. The total mass, angular momentum and entropy are all affected by the factor $\sqrt{1+s}$, arising due to LV effect. Our results differ from those obtained by the conventional thermodynamic method in Ref. [37], where $M=\frac{\mu}{8G}$, $J=\frac{j}{8G}$ and $S=\frac{\sqrt{1+s}\pi r_{{}_{\mathrm{H}}}}{2G}$, with constant G having been retrieved. On the other hand, the Bekenstein-Hawking entropy (41) was straightforwardly derived as a conserved charge, and the first law is perfectly satisfied. This is an advantage of the SPSM. When $s\rightarrow 0$ , this solution reduces to the exact BTZ black hole solution, and $M=\frac{\mu}{4G}$, $J=\frac{j}{4G}$, $S=\frac{\pi r_{{}_{\mathrm{H}}}}{G}$ (up to factor $\frac{1}{2}$ , which comes from the choice of the gravitational coupling constant $\kappa=4\pi G$ in original black hole solution in Ref. [37] and also in our case). Although the bumblebee field strength $B_{\mu\nu}$ is analogous to the electromagnetic field strength $F_{\mu\nu}$, the $\mathrm{U(1)}$ internal gauge symmetry is broken in the bumblebee theory. As a result, there are no corresponding $\mathrm{U(1)}$ conserved charges (such as ‘electromagnetic charges $Q_{\mathrm{I}}$’) and no terms like $\Phi^{\mathrm{I}}_{\mathrm{H}}\delta Q_{\mathrm{I}}$ appearing in the first law. At this point, it is different from the electromagnetic theories with the $\mathrm{U(1)}$ symmetry. As a consistent check, in the next section we will derive the microscopic entropy of dual CFT by the AdS/CFT correspondence, and find that the Bekenstein-Hawking entropy (41) is in precise agreement with the microscopic entropy of the dual CFT.
5 Asymptotic symmetries
@@
5.1 AdS/CFT correspondence
Setting $\mu=0=j$ in the metric and bumblebee field of the BTZ-like black hole (4), we can get the $\mathrm{AdS}_{3}$ background spacetime
@@ 5.1 AdS/CFT correspondence
By using the Brown-Henneaux boundary conditions [8]
@@
$i\left[\xi^{+}_{m},\xi^{-}_{n}\right]=$ $0.$ (48)
For the generator $\xi^{+}_{m}$, the Poisson bracket algebra (2) becomes
@@
$i\oint_{S}\bm{k}_{\xi^{-}_{m}}(\delta_{\xi^{-}_{n}}\bar{\Phi},\bar{\Phi})=\frac{1}{12}\frac{3\sqrt{1+s}l}{G}m^{3}\delta_{m+n,0},$ (51)
where $\tilde{L}^{+}_{m}\equiv H_{\xi^{+}_{m}}$, $\tilde{L}^{-}_{m}\equiv H_{\xi^{-}_{m}}$, and $\bar{\Phi}$ is the background (5.1). To contain centerless subalgebra for m=1,0,+1m=-1,0,+1, we shift the zero modes of the charges
@@
By using the Cardy formula [60, 69], and modified by the LV factor $\sqrt{1+s}$, the microscopic entropy of the dual CFT is
@@
The zero modes of $\tilde{L}^{+}_{m}$ and $\tilde{L}^{-}_{n}$ for the BTZ-like black hole (4) are
@@
5.2 Warped conformal symmetries in near horizon region
In this subsection we will use the Detournay-Smoes-Wutte boundary conditions [58] to study the asymptotic symmetries of the near horizon extremal BTZ-like black hole.
In the extremal case, the BTZ-like metric and the bumblebee field (4) are
@@ 5.2 Warped conformal symmetries in near horizon region
By taking limit $\varepsilon\rightarrow 0$, we can obtain the near horizon geometry of the extremal BTZ-like metric and bumblebee field
@@
We now impose the Detournay-Smoes-Wutte boundary conditions in Ref. [58] on the near horizon geometry of the extremal BTZ-like metric and bumblebee field (5.2), which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity [59], and it is given by a finite coordinate transformation
@@
where $\mathcal{H}(\tau)\equiv\mathcal{F^{\prime\prime}}(\tau)/\mathcal{F^{\prime}}(\tau)$. The near horizon geometry (5.2) is restored when $\mathcal{G^{\prime}}(\tau)=0=\mathcal{H}(\tau)$.
@@
$i\left[j_{m},j_{n}\right]=$ $0.$ (66)
To obtain finite and non-zero charges (2), we take $\partial{\Sigma}$ to be the circle S of constant $(r,\varphi)$ and take the limit $r\rightarrow\infty$. By taking variations of the metric and bumblebee field
@@
the corresponding charge variation (2) associated with $\xi$ can be given by
$\delta H_{\xi}=\frac{\sqrt{1+s}l}{16\pi G}\int^{L}_{0}d\tau\left(4\mu\zeta(\tau)\delta\mathcal{G^{\prime}}(\tau)-2\mu\epsilon(\tau)\delta\mathcal{G^{\prime}}(\tau)^{2}+\epsilon^{\prime}(\tau)\delta\mathcal{H}(\tau)+\epsilon(\tau)\delta\mathcal{H}(\tau)^{2}\right),$ (68)
which is integrable. Under the background (5.2), the charges are given by
$\hat{L}_{n}=$ $H_{l_{n}}=\frac{\sqrt{1+s}l}{16\pi G}\int^{L}_{0}d\tau\left(-2\mu\mathcal{G^{\prime}}(\tau)^{2}+\frac{4\pi in}{L}\mathcal{H}(\tau)+\mathcal{H}(\tau)^{2}\right)\frac{L}{2\pi}e^{2\pi in\tau/L},$
$\hat{J}_{n}=$ $H_{j_{n}}=\frac{\sqrt{1+s}l}{16\pi G}\int^{L}_{0}d\tau\left(4\mu\mathcal{G^{\prime}}(\tau)\right)\frac{L}{2\pi i}e^{2\pi in\tau/L}.$ (69)
The Poisson bracket algebra (2) of $\hat{L}_{n}$ and $\hat{J}_{n}$ is a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, the symmetry algebra of a WCFT,
@@
6 Conclusions and Outlooks
In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole in Einstein-bumblebee theory of gravity. We explicitly computed the black hole mass, angular momentum and entropy, demonstrating that these quantities are influenced by LV parameter. Our results differ from those obtained in Ref. [37] using the conventional thermodynamic method. Additionally, we derived the first law of black hole thermodynamics within this framework. As an asymptotical AdS spacetime, we studied the AdS/CFT correspondence and obtained the asymptotic charge algebra, which are two copies of the Virasoro algebra with non-trivial central charges. By applying the Cardy formula, we computed the microscopic entropy of the dual CFT, and found it is in exact agreement with the Bekenstein-Hawking entropy calculated by SPSM. Thus, we confirmed our results for the conserved charges through entropy matching and by verifying the consistency of the first law of black hole thermodynamics. Furthermore, in the near horizon region, we used the Detournay-Smoes-Wutte boundary conditions to study the near horizon geometry of extremal BTZ-like black hole. We obtained a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of WCFT. And the central charge $\hat{c}$ of the Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra is found to be identical to the central charge $c$ of the Virasoro algebra in AdS/CFT correspondence.
Evidently, the conserved charges and central charges are affected by the LV parameter. We speculate that similar modifications arise in other bumblebee black hole solutions [34, 35, 39, 40, 41, 42], which may also affect the thermodynamic properties. We expect this deviation from General Relativity, as well as the LV parameter for four-dimensional bumblebee gravity, can be further studied and constrained by using astrophysical data, including solar system test [34], Event Horizon Telescope (EHT) observables of black hole shadows [46, 47, 48, 49], quasi-periodic oscillations (QPO) [52], and black hole X-ray data [53]. On the other hand, our study provides a first step towards building a holographic dual for near horizon geometry of bumblebee black holes. Through the Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, as well as the corresponding central charges and Kac-Moody level, many properties can be further explored, for example, deriving a Cardy-type formula in WCFT, exploring holographic thermodynamics, studing holographic entanglement entropy, phase transitions, complex, bulk reconstruction, calculating correlation functions and studing anomalies. Furthermore, the enhanced ASG of $\mathrm{WAdS_{3}}$ spacetimes consist of a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which inspires us to search for exact $\mathrm{WAdS_{3}}$ and warped black hole solutions in the bumblebee gravity. These are potential future research directions along this line.
Acknowledgements